OpenStudy (anonymous):

completing the square 8x=4x^2-2 help pls

OpenStudy (anonymous):

4x^2 - 8x - 2 = 0 4x^2 - 8x + 4 - 6 = 0 (2x-2)^2 - 6 = 0 4*(x-1)^2 - 6 = 0 => (x-1)^2 = 3/2 => x-1 = +/- SQRT(3/2)